What is the minimum concentration of Cl- necessary to begin precipitation of AgCl(s) (Ksp = 1.77×10-1) in a solution of 5.0×10-5M Ag+?
Answer #1

Ag+ + Cl- = AgCl

Ksp (AgCl) = [Ag+] [Cl-]

given Ksp (AgCl) = 1.77×10-1

[Ag+] = 5.0×10-5 M

[Cl-] = Ksp / [Ag+]

[Cl-] = 1.77×10-1 / 5.0×10-5

[Cl-] = 3540 M or 3.5 x 103 M